# The role of magnetic fields in shaping $\gamma$-ray emission from the Fermi bubbles

### v1

@@ 5 Discussion
Our MHD and CR transport simulations allow us to calculate the fraction of the CR energy which subsequently goes on to power the Fermi bubble emission in $\gamma$-rays, a quantity known as the calorimetric fraction. For the CR luminosity used in this paper (see section 4.2) a pure calorimetric exchange, i.e., 100% of the CR luminosity is radiated as $\gamma$-rays, would correspond to $L_{\gamma}=\frac{1}{3}L_{\mathrm{CR}}=4.3\times 10^{39}$ erg s-1, where the factor 1/31/31 / 3 accounts for the $\gamma$-ray emission produced by $\pi^{0}$ (Ackermann et al., 2012; Wang & Fields, 2018). For our model, $L_{\gamma}=2.7\times 10^{37}$ erg s-1 (see section 4.2), corresponding to a calorimetric fraction of $\sim 6\times 10^{-3}$. This implies that most of the CR energy is not radiated away in $\gamma$-rays from the Fermi bubble region, indicating then that these bubbles are non-calorimetric. However, it should be noted that the enhanced inner Galactic density, at least one order larger than the Galactic halo density at a Galactic radius of 300 pc, has not been included in our simulations. This larger density distribution in the central region would result in the production of further $\gamma$-ray emission from the central region, increasing the calorimetric fraction estimate.
@@
A further simplification used for our results is the assumption of a single temperature of the outflow gas. The outflow gas velocities obtained from our MHD simulations are somewhat smaller than the value of $\sim$300 km s-1 (at the critical radius), motivated by observations (Lockman et al., 2020). For the hot Galactic halo gas, observations have indicated the presence of gas in an even hotter phase (Das et al., 2019b, a; Das et al., 2021; Gupta et al., 2021; Ramesh et al., 2023). A temperature profile, expressed through a multi-thermal phase model, may therefore be warranted to consider in the future. Additionally, further studies of the magnetic field distribution in the halo seem also warranted. The presence of a stronger azimuthal magnetic field than that considered here could lead to an increase in the outflow velocity at the critical radius.
@@ 6 Conclusions and outlook
The Fermi bubbles have frequently been associated with a Galactic outflow. Following our previous work in which a hydrodynamic Galactic breeze (subsonic) model was employed to describe the Fermi bubbles (Tourmente et al., 2023), we have here extended this description using MHD simulations. For our fiducial model, an initial radial magnetic field at the inner radial boundary of 20 $\mu$G and an azimuthal magnetic field of 2 $\mu$G are adopted. This MHD simulation setup leads to a velocity profile largely similar to that obtained previously for our hydrodynamic model. Both the velocity and magnetic field profiles from our MHD simulation of the Galactic halo have subsequently been used as inputs for our CR propagation simulation. Our transport of CRs in the Galactic halo, via both advection and anisotropic diffusion, gave rise to $\gamma$-ray emission through inelastic CR collisions with the low-density hot Galactic halo gas.
@@ Acknowledgements
OT and AMT acknowledge support from DESY (Zeuthen, Germany), a member of the Helmholtz Association HGF. DRL would like to acknowledge that this publication has emanated from research conducted with the financial support of Taighde Éireann – Research Ireland under Grant number 21/PATH-S/9339.
@@ Appendix A Effects of a magnetic field distribution on a thermally driven outflow
An introduction to the topic of a hydrodynamical spherically symmetric outflow can be found in chapter three of Lamers & Cassinelli (1999) based on the work of Parker (1958, 1965). We focus only here on a description for an MHD case.
@@
$\frac{1}{v_{r}}\frac{dv_{r}}{dr}=\frac{1}{r}\left(\frac{2c^{2}_{s}-r\frac{d\Phi_{\mathrm{tot}}}{dr}-v^{2}_{\phi}-\frac{B_{\phi}}{4\pi\rho}\frac{d}{dr}\left(rB_{\phi}\right)}{v^{2}_{r}-c^{2}_{s}}\right).$ (11)
@@
$r_{c}=\frac{r\frac{d\Phi_{\mathrm{tot}}}{dr}}{2c^{2}_{s}+v_{\phi}^{2}-\frac{B^{2}_{\phi}}{4\pi\rho}\left(1-\beta\right)}\rightarrow\begin{cases}\beta=1&\Rightarrow B^{2}_{\phi}\left(1-\beta\right)=0\\ \beta>1&\Rightarrow B^{2}_{\phi}\left(1-\beta\right)<0\\ \beta<1&\Rightarrow B^{2}_{\phi}\left(1-\beta\right)>0.\end{cases}$
For $\beta=1$, the magnetic pressure and the magnetic tension cancel each other, and the position of the critical radius is not modified by the presence of the magnetic field. For $\beta>1$, the magnetic pressure is more important than the magnetic tension, and $r_{\rm c}$ is moved closer to the Galactic centre. For $\beta<1$, the magnetic pressure is less important than magnetic tension, and $r_{\rm c}$ moves farther from the Galactic centre. This last case allows for increasing the thermal pressure for keeping a critical radius at 1 kpc. This competition between the influence of the magnetic pressure and the magnetic tension has been previously explored to deduce its influence on the rotation of the Galactic disc (Sánchez-Salcedo & Santillán, 2013; Elstner et al., 2014).
@@ Appendix B Analytic expression for $\mathbf{B_{\phi}(r)}$
[…]
@@
From this expression, one can see that it is not necessary to have $B_{\phi_{0}}\neq 0$ for $B_{\phi}(r)$ to exist. If $B_{\phi_{0}}=0$, $v_{\phi_{0}}$ and $B_{r_{0}}$ must be different from zero. Moreover, the combination $B_{r_{0}}v_{\phi_{0}}$ decreases $\propto r^{-2-1+\alpha}$. A large value $B_{r_{0}}$ or $v_{\phi_{0}}$ is then necessary for $B_{\phi}(r)$ to reach a few $\mu$G $1~\mathrm{kpc}\leq z\leq 10~\mathrm{kpc}$. Also, a too large value for $v_{\phi_{0}}$ would disrupt the propagation of the outflow, as it would compete with $v_{r_{0}}$. It must also be noted that the distribution of $B_{\phi}(r)$ depends mostly on the index $\alpha$ representing the power-law distribution for $v_{r}(r)$. For the Galactic breeze profile (see Fig. 2) for $z>1$ kpc, $\alpha\approx 1$. The distribution of $B_{\phi}(r)$ is then not steep and looks quasi-constant along $r$.
Appendix C The expression of the diffusion coefficient
@@
[…]
Appendix D Energy flux distribution for latitude between 30 and 40
@@
Figure 7: Energy flux distribution along the Galactic longitude for a Galactic latitude between $40^{\circ}$ and $50^{\circ}$. The line styles, shaded regions and normalisation are identical to Fig. 6.
Appendix E Gamma-ray energy flux distribution for an isotropic diffusion model
@@
Figure 8: Spatial CR density distribution in the Galactic halo for an energy range of $E_{\mathrm{CR}}=10~-~30$ GeV for the isotropic diffusion model. The colours and contour lines are the same as for Fig. 4
Figure 9: $\gamma$-ray emission map for an energy range of 1 to 3 GeV obtained from an isotropic diffusion of CRs. A Mollweide projection has been used. The colour bar and mask are similar to Fig. 5. The $\gamma$-ray emission has been produced by $pp$ collisions following the advection of CRs into the hydrostatic density distribution of the Galactic halo. The resulting emission produces a bubble shape for one half hemisphere that is larger than for an anisotropic CR diffusion (see Fig. 5).
Figure 10: $\gamma$-ray energy flux distribution as a function of Galactic longitude, $l$, for a Galactic latitude between $40^{\circ}$ and $50^{\circ}$, assuming isotropic CR diffusion. The line styles and shaded regions are the same as for Figs. 6 and 7.

### v2

@@ 5 Discussion
Our MHD and CR transport simulations allow us to calculate the fraction of the CR energy which subsequently goes on to power the Fermi bubble emission in $\gamma$-rays, a quantity known as the calorimetric fraction. For the CR luminosity used in this paper (see section 4.2) a pure calorimetric exchange, i.e., 100% of the CR luminosity is radiated as $\gamma$-rays, would correspond to $L_{\gamma}=\frac{1}{3}L_{\mathrm{CR}}=4.3\times 10^{39}$ erg s-1, where the factor $1/3$ accounts for the $\gamma$-ray emission produced by $\pi^{0}$ (Ackermann et al., 2012; Wang and Fields, 2018). For our model, $L_{\gamma}=2.7\times 10^{37}$ erg s-1 (see section 4.2), corresponding to a calorimetric fraction of $\sim 6\times 10^{-3}$. This implies that most of the CR energy is not radiated away in $\gamma$-rays from the Fermi bubble region, indicating then that these bubbles are non-calorimetric. However, it should be noted that the enhanced inner Galactic density, at least one order larger than the Galactic halo density at a Galactic radius of 300 pc, has not been included in our simulations. This larger density distribution in the central region would result in the production of further $\gamma$-ray emission from the central region, increasing the calorimetric fraction estimate.
@@
A further simplification used for our results is the assumption of a single temperature of the outflow gas. The outflow gas velocities obtained from our MHD simulations are somewhat smaller than the value of $\sim$300 km s-1 (at the critical radius), motivated by observations (Lockman et al., 2020). For the hot Galactic halo gas, observations have indicated the presence of gas in an even hotter phase (Das et al., 2019a, b, 2021; Gupta et al., 2021; Ramesh et al., 2023). A temperature profile, expressed through a multi-thermal phase model, may therefore be warranted to consider in the future. Additionally, further studies of the magnetic field distribution in the halo seem also warranted. The presence of a stronger azimuthal magnetic field than that considered here could lead to an increase in the outflow velocity at the critical radius.
@@ 6 Conclusions and outlook
[…]
@@ Acknowledgements
The authors would like to thank the anonymous reviewer for their valuable comments which helped to improve the manuscript. OT and AMT acknowledge support from DESY (Zeuthen, Germany), a member of the Helmholtz Association HGF. DRL would like to acknowledge that this publication has emanated from research conducted with the financial support of Taighde Éireann – Research Ireland under Grant number 21/PATH-S/9339.
@@ Appendix A Effects of a magnetic field distribution on a thermally driven outflow
An introduction to the topic of a hydrodynamical spherically symmetric outflow can be found in chapter three of Lamers and Cassinelli (1999) based on the work of Parker (1958, 1965). We focus only here on a description for an MHD case.
@@
$\frac{1}{v_{r}}\frac{dv_{r}}{dr}=\frac{1}{r}\left(\frac{2c^{2}_{s}-r\frac{d\Phi_{\mathrm{tot}}}{dr}+v^{2}_{\phi}-\frac{B_{\phi}}{4\pi\rho}\frac{d}{dr}\left(rB_{\phi}\right)}{v^{2}_{r}-c^{2}_{s}}\right).$ (14)
@@
$r_{c}=\frac{2c^{2}_{s}+v_{\phi}^{2}-\frac{B^{2}_{\phi}}{4\pi\rho}\left(1-\beta\right)}{\frac{d\Phi_{\mathrm{tot}}}{dr}}\rightarrow\begin{cases}\beta=1&\Rightarrow B^{2}_{\phi}\left(1-\beta\right)=0\\ \beta>1&\Rightarrow B^{2}_{\phi}\left(1-\beta\right)<0\\ \beta<1&\Rightarrow B^{2}_{\phi}\left(1-\beta\right)>0.\end{cases}$
For $\beta=1$, the magnetic pressure and the magnetic tension cancel each other, and the position of the critical radius is not modified by the presence of the magnetic field. For $\beta>1$, the magnetic pressure is more important than the magnetic tension, and $r_{\rm c}$ is moved closer to the Galactic centre. For $\beta<1$, the magnetic pressure is less important than magnetic tension, and $r_{\rm c}$ moves farther from the Galactic centre. This last case allows for increasing the thermal pressure for keeping a critical radius at 1 kpc. This competition between the influence of the magnetic pressure and the magnetic tension has been previously explored to deduce its influence on the rotation of the Galactic disc (Sánchez-Salcedo and Santillán, 2013; Elstner et al., 2014).
@@ Appendix B Analytic expression for $\mathbf{B_{\phi}(r)}$
$\nabla\times\left(\textbf{{v}}\times\textbf{B}\right)=0.$
@@
[…]
Appendix C Effect of compression and advection terms for the induction equation
The induction equation (Eq. 3) can be re-expressed as
$\frac{\partial\mbox{$B$}}{\partial t}=-\mbox{$B$}\left(\nabla\cdot\mbox{$v$}\right)-\left(\mbox{$v$}\cdot\nabla\right)\mbox{$B$}+\left(\mbox{$B$}\cdot\nabla\right)\mbox{$v$},$ (15)
where the terms on the right-hand side correspond to compression, advection, and stretching of the magnetic field, respectively. Fig. 7 presents the distribution of the different terms corresponding to the compression (in red), the advection (in green) and the total of those terms (in orange). The solid lines correspond to the radial term and dotted lines correspond to the azimuthal term, respectively. The stretching term has been omitted as it is negligible in comparison of the two other terms. The shell is predominantly formed by the azimuthal advection term. It should be noted that the time scale on which $B_{\phi}$ changes, that is $\sim$ Gyr scale, is significantly larger than the dynamical age of the Fermi bubbles, 300 Myr for our simulations, ensuring that the CR distribution reaches a steady-state out of a distance of 10 kpc.
Figure 7: Distribution of the induction equation components (see Eq. 15) for the last time step, corresponding to compression (in red) and advection (in green) and the total of those two terms (in orange) along $r$. Solid lines correspond to the radial term and dotted lines correspond to the azimuthal term.
Appendix D The expression of the diffusion coefficient
@@
$\mbox{$D$}=D_{\perp}\left(\mbox{$I$}-\frac{\mbox{$B$}}{|\mbox{$B$}|}\frac{\mbox{$B$}}{|\mbox{$B$}|}\right)+D_{\parallel}\frac{\mbox{$B$}}{|\mbox{$B$}|}\frac{\mbox{$B$}}{|\mbox{$B$}|}.$
Appendix E Energy flux distribution for latitude between 30 and 40
@@
Figure 8: Energy flux distribution along the Galactic longitude for a Galactic latitude between $30^{\circ}$ and $40^{\circ}$. The line styles, shaded regions and normalisation are identical to Fig. 6.
Appendix F Gamma-ray energy flux distribution for an isotropic diffusion model
@@
Figure 9: Spatial CR density distribution in the Galactic halo for an energy range of $E_{\mathrm{CR}}=10~-~30$ GeV for the isotropic diffusion model. The colours and contour lines are the same as for Fig. 4.
Figure 10: $\gamma$-ray emission map for an energy range of 1 to 3 GeV obtained from an isotropic diffusion of CRs. A Mollweide projection has been used. The colour bar and mask are similar to Fig. 5. The $\gamma$-ray emission has been produced by pppp collisions following the advection of CRs into the hydrostatic density distribution of the Galactic halo. The resulting emission produces a bubble shape for one half hemisphere that is larger than for an anisotropic CR diffusion (see Fig. 5).
Figure 11: $\gamma$-ray energy flux distribution as a function of Galactic longitude, $l$, for a Galactic latitude between $40^{\circ}$ and $50^{\circ}$, assuming isotropic CR diffusion. The line styles and shaded regions are the same as for Figs. 6 and 8.